Circular No. 8335

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004bm, 2004bn, 2004bo, 2004bp Four apparent supernovae have been reported from unfiltered CCD

survey images: SN 2004bp by M. Armstrong (cf. IAUC 8320) and the remainder by D. Singer, J. Burket, and W. Li (LOSS/KAIT; IAUC 8329).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004bm Apr. 25.21 10^{h}52^{m}35.^{s}33 +22^{o}56′05′′.5 17.5 6′′.2 W, 2′′.5 N
2004bn Apr. 23.23 10 52 31.06 + 7 13 41.2 18.8 1′′.3 W, 11′′.5 N
2004bo Apr. 30.28 13 26 06.39 −19 46 32.6 17.0 19′′.8 W, 3′′.8 N
2004bp May 3.94 11 09 53.25 +37 00 06.4 18.0 14′′.2 E, 0′′.6 S

Additional approximate magnitudes from the respective discoverers: SN 2004bm in NGC 3437, Mar. 15.22 UT, [19.0; May 2.23, 17.5. SN 2004bn in NGC 3441, Jan. 2.40, [19.5; Apr. 24.19, 18.8; May 4.21, 18.7. SN 2004bo in ESO 576-G54, Mar. 31.32, [19.0; Apr. 13.31, 17.0; May 5.35, 17.0. SN 2004bp in MCG +06-25-14, 1989 Jan. 14, [20.8 (Palomar Sky Survey, red); 1990 Apr. 21, [22.5 (Palomar Sky Survey, blue); 2002 Feb. 22, [19.0; 2004 Apr. 21.99, 19.5 (hint); May 4.943, 18.0.

NOTICE CONCERNING FAINT POSSIBLE SUPERNOVAE Extensive discussions between the Central Bureau and supernova ob-

servers in the last couple of years — largely regarding the problem of very faint, spectroscopically unconfirmed supernova suspects — culminated in the decision, made at the meeting of the IAU Working Group on Supernovae (WGS) at the IAU General Assembly in Sydney last July, to establish an interactive webpage (cohosted by the CBAT and WGS) that will allow registered users to post discovery observations immediately under a provisionaldesignation scheme so that the entire community can have ready access to them in helping to confirm them as supernovae. The interactive webpage is still under construction, but with observers’ eagerness to get their data announced quickly, the Central Bureau has established a preliminary, nonin-

teractive site at URL http://cfa-www.harvard.edu/iau/CBAT PSN.html.

Contributors are asked to read the instructions at that website and submit data in the standard format described therein. Objects listed at this website will have provisional designations tied in to the date of discovery; any objects later confirmed as supernovae will receive formal supernova designations and be announced, as usual, on these Circulars.

2004 May 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green